Please find the avg Molarity and standard deviation. Thank you. And show calculations.
?? Laboratory Experiment . Concentration of Unknown Acid piosnsll) noibuloa bios pios Trial 3 Trial 1 Trial 2 olume of unknown acid Average molarity of NaOH from above mL of NaOH at equivalence point Molarity of unknown acid Average molarity (show calculations) (HOM M GM standard deviation- Sc
Answer #1

Trial 1

M1V1 (acid) = M2V2 (base)

M1 = (M2* V2)/(V1)

(0.9788 * 15.6)/ 25 = 0.6108M

Trial 2

M1V1 (acid) = M2V2 (base)

M1 = (M2* V2)/(V1)

(0.9788 * 15.5)/ 25 = 0.6069 M

Trial 3

M1V1 (acid) = M2V2 (base)

M1 = (M2* V2)/(V1)

(0.9788 * 15.6)/ 25 = 0.6108M

Average molarity = (0.6108 + 0.6069 + 0.6108)/3

= 0.6091

Standard deviation = Tauer ge

= 0.0023